# Explaining Sources of Uncertainty in Automated Fact-Checking

### v1

@@ Abstract
Understanding sources of a model’s uncertainty regarding its predictions is crucial for effective human-AI collaboration. Prior work proposes to use numerical uncertainty or hedges (“I’m not sure, but…”), which do not explain uncertainty arising from conflicting evidence, leaving users unable to resolve disagreements or rely on the output. We introduce CLUE (Conflict-&Agreement-aware Language-model Uncertainty Explanations), the first framework to generate natural language explanations of model uncertainty by: (i) identifying relationships between spans of text that expose claim-evidence or inter-evidence conflicts/agreements driving the model’s predictive uncertainty in an unsupervised way; and (ii) generating explanations via prompting and attention steering to verbalize these critical interactions. Across three language models and two fact-checking datasets, we demonstrate that CLUE generates explanations that are more faithful to model uncertainty and more consistent with fact-checking decisions than prompting for explanation of uncertainty without span-interaction guidance. Human evaluators find our explanations more helpful, more informative, less redundant, and more logically consistent with the input than this prompting baseline. CLUE requires no fine-tuning or architectural changes, making it plug-and-play for any white-box language model. By explicitly linking uncertainty to evidence conflicts, it offers practical support for fact-checking and readily generalizes to other tasks that require reasoning over complex information.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly prevalent in high-stakes tasks that involve reasoning about information reliability, such as fact-checking (Wang et al., 2024; Fontana et al., 2025). To foster effective use of such models in fact-checking tasks, these models must explain the rationale for their predictions (Atanasova et al., 2020; Kotonya and Toni, 2020).
Figure 1: Example of claim and evidence documents, alongside span interactions for uncertainty and generated natural language explanations.
However, current methods in automated fact-checking have been criticised for their failure to address practical explainability needs of fact-checkers (Warren et al., 2025) and for their disconnect from the tasks typically performed by fact-checkers (Schlichtkrull et al., 2023). For example, although fact-checking involves complex reasoning about the reliability of (often conflicting) evidence, existing automatic fact-checking techniques focus only on justifying the verdict (Atanasova et al., 2020; Stammbach and Ash, 2020; Zeng and Gao, 2024). Such methods do not explain the uncertainty associated with their predictions, which is crucial for their users to determine whether some of the uncertainty is resolvable, and if so, which aspects of this uncertainty within the evidence to address (e.g., by retrieving additional information) (Warren et al., 2025).
@@
Uncertainty in model predictions is often communicated through numerical scores (e.g., “I am 73% confident”), however, such metrics can be hard to contextualize and lack actionable insights for end-users (Zimmer, 1983; Wallsten et al., 1993; van der Waa et al., 2020; Liu et al., 2020). Recent efforts have instead used natural language expressions (e.g., “I’m not sure”) to convey uncertainty (Steyvers et al., 2025; Yona et al., 2024; Kim et al., 2024), but such expressions often fail to faithfully reflect model uncertainty (Yona et al., 2024), and users may overestimate model confidence (Steyvers et al., 2025). Existing explainable fact-checking systems exhibit two critical limitations: they focus solely on justifying veracity predictions through generic reasoning summaries of the input sequence (see Figure 2), while neglecting to (1) communicate model uncertainty or (2) explicitly surface evidentiary conflicts and agreements that relate to it. This constitutes a fundamental methodological gap, as effective fact-checking requires precisely identifying the sources of uncertainty, for example from conflicting evidence, to guide targeted verification Graves (2017); Micallef et al. (2022).
@@
We propose CLUE, a pipeline that generates natural language explanations (NLEs) of model uncertainty by explicitly capturing conflicts and agreements in the input (e.g., a claim and its supporting or refuting evidence). The pipeline first identifies the salient span-level interactions that matter to the prediction of the model through an unsupervised approach, providing an input-feature explanation that highlights key relationships between separate input segments (e.g., claim and evidence) Ray Choudhury et al. (2023). These interactions have been shown to be both faithful to the model and plausible to humans Sun et al. (2025). CLUE then converts these signals into uncertainty-aware explanations by explicitly discussing the interactions, the conflict/agreement relations they express and how they contribute to uncertainty regarding the verdict. CLUE does not require gold-label explanations, avoids fine-tuning, and operates entirely at inference time.
@@
Across three language models (§4.2) and two fact-checking datasets (§4.1), we evaluate two variants of CLUE. Automatic metrics show that both variants generate explanations that are more faithful to each model’s uncertainty and agree more closely with the gold fact-checking labels than a prompting baseline that lacks conflict-/agreement-span guidance (§5.5). Human study participants likewise judge CLUE explanations as more helpful, more informative, less redundant, and more logically consistent with the input. We also observe a trade-off between two variants of our CLUE framework, one attains higher faithfulness, the other higher plausibility, highlighting a promising avenue for future work to achieve both simultaneously (§5.5).
@@
Figure 2: Explanations produced by earlier systems, e-FEVER (Stammbach and Ash, 2020), Explain-MT (Atanasova et al., 2020), and JustiLM (Zeng and Gao, 2024), compared with those from our CLUE framework. CLUE is the only approach that explicitly traces model uncertainty to the conflicts and agreements between the claim and multiple evidence passages.
2 Related Work
2.1 Uncertainty Quantification in LLMs
Recent work on LLM uncertainty quantification primarily relies on logit-based methods such as answer distribution entropy Kadavath et al. (2022), summing predictive entropies across generations Malinin and Gales (2021), and applying predictive entropy to multi-answer question-answering Yang et al. (2025). Estimating uncertainty in long-form tasks involves measuring semantic similarity between responses (Duan et al., 2024; Kuhn et al., 2023; Nikitin et al., 2024). Given that logit-based uncertainty quantification is infeasible for closed-source black-box models, alternative approaches have depended on verbalizing confidence directly (Lin et al., 2022; Mielke et al., 2022), though these measures are overconfident and unreliable (Yona et al., 2024; Tanneru et al., 2024). Other approaches measure output diversity across paraphrased prompts (Zhang et al., 2024a; Chen and Mueller, 2024), but this technique can introduce significant computational overhead and conflate model uncertainty with prompt-induced noise, obscuring interpretability. Accordingly, in this work, we focus on the uncertainty of open-source models, which are readily accessible and widely used. We adopt predictive entropy, a straightforward white-box metric computed from the model’s answer logits, as our uncertainty measure for fact-checking tasks. This choice balances interpretability and computational efficiency while avoiding potential noise introduced by multiple prompts.
2.2 Linguistic Expressions of Uncertainty
Numerical uncertainty estimates do not address the sources of uncertainty, and are therefore difficult for end-users, such as fact-checkers, to interpret and act upon (Warren et al., 2025). Linguistic expressions of uncertainty may be more intuitive for people to understand than numerical ones, (Zimmer, 1983; Wallsten et al., 1993; Windschitl and Wells, 1996), and recent work has proposed models that communicate uncertainty through hedging phrases such as “I am sure” or “I doubt” (Mielke et al., 2022; Lin et al., 2022; Zhou et al., 2023; Tian et al., 2023; Xiong et al., 2023; Ji et al., 2025; Zheng et al., 2023; Farquhar et al., 2024). However, these expressions are not necessarily faithful reflections of the model’s uncertainty (Yona et al., 2024) and tend to overestimate the model’s confidence  (Tanneru et al., 2024), risking misleading users (Steyvers et al., 2025). Moreover, they do not explain why the model is uncertain. In this paper, we propose a method that explains sources of model uncertainty by referring to specific conflicting or concordant parts of the input that contribute to the model’s confidence in the output. This approach ensures a more faithful reflection of model uncertainty and provides users with a more intuitive and actionable understanding of model confidence.
2.3 Generating Natural Language Explanations for Fact-Checking
Natural language explanations provide justifications for model predictions designed to be understood by laypeople (Wei Jie et al., 2024). NLEs have typically been evaluated by measuring the similarity between generated NLEs and human-written reference explanations using surface-level metrics such as ROUGE-1 (Lin, 2004) and BLEU (Papineni et al., 2002). In fact-checking, supervised methods have been proposed that involve extracting key sentences from existing fact-checking articles and using them as explanations (Atanasova et al., 2020). Later work proposed a post-editing mechanism to enhance the coherence and fluency of explanations, (Jolly et al., 2022), while others have fine-tuned models on data collected from fact-checking websites to generate explanations (Feher et al., 2025; Raffel et al., 2020; Beltagy et al., 2020). Recent work has shifted towards few-shot methods requiring no fine-tuning, for example, using few-shot prompting with GPT-3  (Brown et al., 2020) to produce evidence summaries as explanations (Stammbach and Ash, 2020) incorporating a planning step before explanation generation (Zhao et al., 2024) to outperform standard prompting approaches, and generating fact-checking justifications based on retrieval-augmented language models (Zeng and Gao, 2024). However, existing methods are often not faithful to model reasoning (Atanasova et al., 2023; Siegel et al., 2024, 2025), have limited utility in fact-checking Schmitt et al. (2024), and fail to address model uncertainty, which has been identified as a key criterion for fact-checking Warren et al. (2025).
To this end, we introduce the first framework designed for the task of explaining sources of uncertainty in multi-evidence fact-checking. Our method analyzes span-level agreements and conflicts correlated with uncertainty scores. Unlike conventional approaches that aim to replicate human NLEs (prioritising fluency or plausibility over faithfulness to model reasoning), our method generates explanations that are both faithful to model uncertainty and helpful to people in a fact-checking context.
@@ 3.1 Preliminaries and Overall Framework
Each input instance is a triple $X=(C,E_{1},E_{2})$ consisting of a claim $C$ and two evidence pieces $E_{1},E_{2}$. Note that, in this work, we set the number of evidence pieces to two for simplicity. For clarity, we denote their concatenation as $X=[x_{1},\dots,x_{|C|+|E_{1}|+|E_{2}|}]$. The task label comes from the set $\mathcal{Y}=\{\textsc{Supports},\textsc{Refutes},\textsc{Neutral}\}$.
@@
    Explanation generation. The model receives the extracted spans as soft constraints and produces a natural-language rationale $Y_{R}=[y_{1}^{\prime},\dots,y_{r}^{\prime}]$ along with its predicted label $\hat{y}$ to the identified interactions (§ 3.4).
@@ 3.2 Predictive Uncertainty Score Generation
To quantify model uncertainty for generating an answer label on a specific input sequence, we follow previous work and calculate predictive uncertainty with entropy theory, which does not require multiple runs and is widely used in open-source models.
@@ 3.3 Conflict and Agreement Span Interaction Identification for Answer Uncertainty
For each ordered input part pair $(F,T)\in\{(C,E_{1}),(C,E_{2}),(E_{1},E_{2})\}$, we follow previous work Ray Choudhury et al. (2023); Sun et al. (2025) to extract the important span interactions and their importance score to model’s answer by (i) identifying the most important attention head to the model’s answer prediction from its final layer, (ii) obtaining its attention matrix $\mathbf{A}\in\mathbb{R}^{(|F|+|T|)\times(|F|+|T|)}$, and (iii) symmetrizing the cross-part scores:
@@
Treating $a^{\prime}_{p,q}$ as edge weights yields a bipartite token graph, which we partition into contiguous spans with the Louvain algorithm (Blondel et al., 2008). Given a $\text{span}_{w}\subset F$ and a $\text{span}_{v}\subset T$, their interaction importance is
@@
The scored interactions for $(S,T)$ form $S_{(S,T)}=\{((\text{span}_{w},\text{span}_{v}),\,a_{wv})\}$.
@@
To tag each span pair as an agreement, disagreement, or unrelated, we prompt GPT-4o (OpenAI Team, 2024) [^1] https://openai.com/index/hello-gpt-4o/ to assign a label $r_{wv}\in\{\text{agree},\text{disagree},\text{unrelated}\}$,balancing scalability and accuracy (See templates in App. H.6).
After labeling all three pairs, the complete interaction set for instance $X$ is
$S_{R}=S_{R}(C,E_{1})\;\cup\;S_{R}(C,E_{2})\;\cup\;S_{R}(E_{1},E_{2}),$ (4)
where, for example, $S_{R}(C,E_{1})=\{((\text{span}_{w},\text{span}_{v}),\,a_{wv},\,r_{wv})\}$. Each element links two spans with an importance score and a relation label, thereby supplying the conflict- or agreement-span interactions used in later stages.
@@ 3.4 Uncertainty Natural Language Explanation Generation
To convert the extracted conflict- and agreement spans to rationales for model uncertainty, we rely on two complementary mechanisms. (i) Instruction-driven prompting embeds the spans directly in the input so the model is instructed which segments to reference. (ii) Intrinsic attention steering guides the model’s own attention toward those same segments while it is generating the rationale. Both mechanisms use self-rationalization: the model first states its verdict $\hat{y}$ and then explains $Y_{R}$, a sequencing shown to improve faithfulness over pipeline approaches  (Wiegreffe et al., 2021; Marasovic et al., 2022; Siegel et al., 2025).
Instruction-based NLE.
For each instance $X$, we rank all labelled interactions by importance and keep the top $K=3$, denoted $S_{R}^{(K)}$, to avoid overly long explanations. These three span pairs are slotted into a three-shot prompt (See App.F.1), which instructs the model to explain how the highlighted agreements or conflicts influence its confidence. Finally, the standard transformer decoding process outputs both the predicted label $\hat{y}$ and the accompanying explanation $Y_{R}$.
@@
Instead of explicit instructions, we can guide generation by modifying attention on the fly with PASTA (Zhang et al., 2024b). Starting from the same $S_{R}^{(K)}$, we collect all token indices that fall inside any selected span,
=$\mathcal{I}=\bigl{\{}p\;:\;(\text{span}_{w},\text{span}_{v})\!\in\!S_{R}^{(K)},\;p\!\in\!\text{span}_{w}\cup\text{span}_{v}\bigr{\}}.$ (5)
@@ 4.2 Models
We compare three generation strategies for NLEs towards model uncertainty:
  • PromptBaseline: A three-shot prompt baseline extending prior few-shot NLE work (Stammbach and Ash, 2020; Zeng and Gao, 2024; Zhao et al., 2024) by explicitly asking the model to highlight conflicting or supporting spans that shape its uncertainty (See prompt template in App. F.1).
  • CLUE-Span: The instruction-based variant of our CLUE method where the extracted span interactions are filled into a three-shot prompt to guide the explanation generation ( §3.4; prompt template in App.F.2).
  • CLUE-Span+Steering: The attention steering variant of our CLUE method in which the same prompt as CLUE-Span is used. Additional attention steering is applied to instinctively guide the model’s explanation generation toward the identified spans ( §3.4; prompt template in App. F.2).
Experiments are run on three recent, open-weight, instruction-tuned LLMs of comparable size: Qwen2.5-14B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct (Qwen Team, 2024), Gemma-2 9B-IT [^4] https://huggingface.co/google/gemma-2-9b-it (Gemma Team, 2024), and OLMo-2-1124-13B-Instruct [^5] https://huggingface.co/allenai/OLMo-2-1124-13B-Instruct (Team OLMo et al., 2024). Each backbone is used consistently across our pipeline for span-interaction extraction, answer prediction, and NLE generation on four NVIDIA A100-SXMS-40GB GPUs. We selected these models to balance capability (reasoning and instruction-following quality) with practical constraints on inference latency and GPU memory.

### v2

@@ Abstract
Human-AI collaboration in knowledge-intensive tasks such as fact-checking requires understanding model uncertainty in multi‑document reasoning amid conflicting/agreeing evidence. Yet, existing methods only express uncertainty as numbers or hedges without revealing which evidence conflicts cause the uncertainty, leaving users unable to resolve disagreements. We present CLUE (Conflict-&Agreement-aware Language-model Uncertainty Explanations), a plug‑and‑play white‑box framework that, to our knowledge, is the first to generate natural‑language explanations of model uncertainty grounded in conflicting/agreeing evidence. CLUE (i) identifies span‑level claim-evidence and inter‑evidence relations that signal conflict or agreement without supervision, and (ii) uses these relations to steer explanation generation, articulating how they drive the model’s uncertainty. Across three language models and two fact‑checking datasets, CLUE produces explanations that more faithfully track model uncertainty and better align with the model’s fact‑checking decisions than span‑agnostic explanation prompting; human raters also judge them more helpful, more informative, less redundant, and more logically consistent with the input. By explicitly tying uncertainty to evidence conflicts and agreements, CLUE supports practical fact‑checking and other tasks that require reasoning over complex, conflicting information.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly prevalent in supporting high-stakes tasks that involve reasoning about information reliability, such as fact-checking (Wang et al., 2024; Fontana et al., 2025). To foster effective use of such models in fact-checking tasks, these models must explain the rationale for their predictions (Atanasova et al., 2020; Kotonya and Toni, 2020).
Figure 1: An example of CLUE output for a claim with two evidence passages: model verdict and certainty, and an uncertainty explanation grounded in conflict/agreement span interactions.
@@
However, current methods in automated fact-checking have been criticised for misaligning with the practical needs of fact-checkers, who must assess source reliability and reconcile conflicting evidence, rather than merely predicting a verdict label (Warren et al., 2025; Schlichtkrull et al., 2023). Current approaches often produce explanations to justify the verdict prediction  (Atanasova et al., 2020; Stammbach and Ash, 2020; Zeng and Gao, 2024), however, they provide no support for communicating model uncertainty or surfacing agreements and conflicts within the evidence (See Fig. 2). Methods that do express uncertainty do so via numeric scores (e.g., “I am 73% confident”) that are hard to contextualise (Zimmer, 1983; Wallsten et al., 1993; van der Waa et al., 2020; Liu et al., 2020) or via natural language expressions (e.g., “I’m not sure”) that often fail to faithfully reflect model uncertainty and can inflate perceived confidence (Steyvers et al., 2025; Yona et al., 2024; Kim et al., 2024). Therefore, existing explainable fact-checking systems usually exhibit two limitations: (1) neglecting to communicate model uncertainty with explanations and (2) failing to surface evidentiary conflicts and agreements that contribute to model uncertainty. This constitutes a fundamental methodological gap, as effective fact-checking requires precisely identifying the sources of uncertainty, for example, from conflicting evidence, to guide targeted verification Graves (2017); Micallef et al. (2022).
@@
To bridge this gap, we propose CLUE, a pipeline that generates natural language explanations (NLEs) of model uncertainty by explicitly capturing conflicts and agreements in the input (e.g., a claim and its supporting or refuting evidence), see Fig. 1. The pipeline first identifies the salient span-level interactions that are essential to the prediction of the model through an unsupervised approach, providing an input-feature explanation that highlights key relationships between separate input segments, e.g., claim and evidence Ray Choudhury et al. (2023). Previous work has shown that these interactions are faithful to the model and plausible to humans Sun et al. (2025, 2026). CLUE then converts these signals into uncertainty-aware explanations by explicitly referring to the interactions, the conflict/agreement relations they express, and how they contribute to uncertainty regarding the verdict. CLUE does not require gold-label explanations or fine-tuning, and operates entirely at inference time.
@@
Across three language models (§3.2) and two fact-checking datasets (§3.1), we evaluate two variants of CLUE, which leverage distinct methods to guide NLE generation based on the extracted key conflicts/agreements (§2.4). Automatic evaluation shows that both variants generate explanations that are more faithful to each model’s uncertainty and agree more closely with the gold fact-checking labels than a prompting baseline that lacks conflict-/agreement-span guidance (§4.5). Human study participants likewise judge CLUE explanations as more helpful, more informative, less redundant, and more logically consistent with the input. We also observe a trade-off between two variants of our CLUE framework, one attains higher faithfulness, the other higher plausibility, highlighting a promising avenue for future work to achieve both simultaneously (§4.5).
@@
Figure 2: Illustrative comparison of verdict-oriented fact-checking explanations (e-FEVER (Stammbach and Ash, 2020), Explain-MT (Atanasova et al., 2020), JustiLM (Zeng and Gao, 2024)) and CLUE’s uncertainty-oriented explanations. In this comparison, CLUE is the only approach that explicitly traces predictive uncertainty to claim-evidence and inter-evidence conflict/agreement across multiple evidence passages.
@@ 2.1 Preliminaries and Overall Framework
Each input instance is a triple $X=(C,E_{1},E_{2})$ consisting of a claim $C$ and two evidence pieces $E_{1},E_{2}$. In this work, we explore the scenarios in which one claim and two or three evidence pieces are presented. [^1] As our framework is based on the pairwise input part interaction, it is readily applicable to scenarios where more claims and evidence pieces are involved. For clarity, we denote their concatenation as $X=[x_{1},\dots,x_{|C|+|E_{1}|+|E_{2}|}]$. The task label comes from the set $\mathcal{Y}=\{\textsc{Supports},\textsc{Refutes},\textsc{Neutral}\}$.
@@
    Explanation generation. The model receives the extracted spans as soft constraints and produces a natural language explanation $Y_{R}=[y_{1}^{\prime},\dots,y_{r}^{\prime}]$ along with its predicted label $\hat{y}$ to the identified interactions (§ 2.4).
@@ 2.2 Predictive Uncertainty Score Generation
To quantify model uncertainty for generating an answer label on a specific input sequence, we follow previous work (Kadavath et al., 2022; Yang et al., 2025) and calculate predictive uncertainty with entropy theory, which does not require multiple runs and is widely used in open-source models.
@@ 2.3 Conflict and Agreement Span Interaction Identification for Answer Uncertainty
For each ordered input part pair $(F,T)\in\{(C,E_{1}),(C,E_{2}),(E_{1},E_{2})\}$, we follow previous work Ray Choudhury et al. (2023); Sun et al. (2025) to extract the important span interactions and their importance score to model’s answer by (i) identifying the most important attention head to the model’s answer prediction from its final layer, (ii) obtaining its attention matrix $\mathbf{A}\in\mathbb{R}^{(|F|+|T|)\times(|F|+|T|)}$, and (iii) symmetrizing the cross-part score $a^{\prime}_{p,q}$ for each possible token interaction $(x_{p},x_{q})$ within $(F,T)$:
@@
Based on these token interactions, we treat each $a^{\prime}_{p,q}$ as an edge weight in a bipartite token graph between the two input parts $(F,T)$. Then the Louvain algorithm (Blondel et al., 2008) is applied to search for communities of tokens with dense intra-cluster and sparse inter-cluster relationships. From each detected token community, one span interaction, $(\text{span}_{w},\text{span}_{v})$, can be extracted by treating the neighboring tokens of the same input part, $F$ and $T$, respectively, as a span. By averaging the token importance score $a^{\prime}_{p,q}$ of each token interaction within $(\text{span}_{w},\text{span}_{v})$, its importance score $a_{wv}$ can be obtained by
@@
The scored interactions for $(F,T)$ form $S^{(F,T)}=\{((\text{span}_{w},\text{span}_{v}),\,a_{wv})\}$.
@@
To tag each span interaction as agreement, disagreement, or unrelated, we use a relation labeler $L_{\text{rel}}$ that can be instantiated with either an open-weight or closed LLM. $L_{\text{rel}}$ takes the full $(C,E_{1},E_{2})$ context together with a candidate span pair and outputs one label in $\{\textit{agree},\textit{disagree},\textit{unrelated}\}$. We implement this using an instruction prompt (see App. C.1) and report the labeling model in Sec. 3.2.
After labelling all three pairs, the complete interaction set for instance $X$ is
$S_{R}=S_{R}^{(C,E_{1})}\;\cup\;S_{R}^{(C,E_{2})}\;\cup\;S_{R}^{(E_{1},E_{2})},$ (4)
where, for example, $S_{R}^{(C,E_{1})}=\{((\text{span}_{w},\text{span}_{v}),\,a_{wv},\,r_{wv})\}$. Each element links two spans with an importance score and a relation label, thereby supplying the conflict- or agreement-span interactions used in later stages.
@@ 2.4 Uncertainty Natural Language Explanation Generation
To convert the extracted conflict- and agreement-span interactions into explanations for model uncertainty, we rely on two complementary mechanisms. (i) Instruction-driven prompting embeds the spans directly in the input so the model is instructed which segments to reference. (ii) Intrinsic attention steering guides the model’s own attention toward those same segments while it is generating the NLE. Both mechanisms use self-rationalization: the model first states its verdict $\hat{y}$ and then explains $Y_{R}$, a sequencing shown to improve faithfulness over pipeline approaches (Wiegreffe et al., 2021; Marasovic et al., 2022; Siegel et al., 2025).
Instruction-driven NLE.
For each instance $X$, we rank all labelled interactions by descending importance scores and keep the top $k=3$, denoted $S_{R}(k)$, to avoid overly long explanations. These three span pairs are slotted into a three-shot prompt (See App. G.1), which instructs the model to explain how the highlighted agreements or conflicts influence its confidence. Finally, the standard transformer decoding process outputs both the predicted label $\hat{y}$ and the accompanying explanation $Y_{R}$.
@@
Instead of explicit instructions, we can guide NLE generation by modifying attention on the fly following the attention steering method in PASTA (Zhang et al., 2024b). Starting from the same $S_{R}(k)$, we collect all token indices that fall inside any selected span,
$\mathcal{I}=\bigl\{p\;:\;(\text{span}_{w},\text{span}_{v})\!\in\!S_{R}(k),\;p\!\in\!\text{span}_{w}\cup\text{span}_{v}\bigr\}.$ (5)
@@ 3.2 Models
We compare three generation strategies for NLEs of model uncertainty:
  • PromptBaseline: A three-shot prompt baseline extending prior few-shot NLE work (Stammbach and Ash, 2020; Zeng and Gao, 2024; Zhao et al., 2024) by explicitly asking the model to highlight conflicting or supporting spans that shape its uncertainty (see prompt template in App. G.1).
  • CLUE-Span: The instruction-driven variant of our CLUE method where the extracted span interactions are filled into a three-shot prompt to guide the explanation generation (§2.4; App. G.2).
  • CLUE-Span+Steering: The attention steering variant of our CLUE method in which the same prompt as CLUE-Span is used. Additional attention steering is applied to instinctively guide the model’s explanation generation toward the identified spans (§2.4; App. G.2).
We run the main experiments on multiple recent open-weight LLMs of comparable scale: Qwen2.5-14B-Instruct (Qwen Team, 2024), Gemma-2-9B-IT (Gemma Team, 2024), and OLMo-2-1124-13B-Instruct (Team OLMo et al., 2025) and DeepSeek-R1-Distill-Qwen-14B (DeepSeek-AI, ) (model cards in App. D). Each backbone is run through CLUE on a single NVIDIA A100-SXM 80GB GPU; we choose these models to balance instruction-following/reasoning ability with inference efficiency (see cost report in App. E).
For span-interaction relation labeling, we instantiate $L_{\text{rel}}$ with GPT-4o (OpenAI Team, 2024) to reduce relation-label noise (see App. C.3 for error propagation analysis). This step is optional: replacing $L_{\text{rel}}$ with the open-weight Qwen2.5-14B-Instruct yields comparable downstream NLE quality with a small drop (App. C.2); we discuss this modularity and future directions in Limitations.